# Unsupervised detection of semantic correlations in big data

### v2

@@ Abstract
Correlations among the features defining a data point force a data set to live in a manifold $\mathcal{M}$ whose geometry encodes the correlation structure of data. The most fundamental property of $\mathcal{M}$ is its intrinsic dimensionBennett [1969] (ID) which is qualitatively defined by the number of independent variables needed to parameterize it without significant information loss.
If one was able to estimate the intrinsic dimension in a generic big data scenario, one would be able to quantify the strength of correlations in real-world datasets, like a corpus of sentences or a large set of measurements from a camera. However, real-world data are represented in feature spaces that are outrageously large, and most of the available methods meet serious problems if the intrinsic dimension of the manifold is of the order of hundreds or moreCamastra and Staiano [2016]. We developed an intrinsic dimension estimator that generalizes the concept of the intrinsic dimension to the case of binary variables. We show that this allows circumventing the curse of dimensionality, estimating reliably the intrinsic dimension even when it is of order 100,000. This result paves the way to a plethora of applications, allowing to address questions which have thus far been, to the best of our knowledge, intractable. For example, we can measure the range of semantic correlations in a corpus of sentences with hundreds of words, and we demonstrate that these correlations are destroyed when shorter, unrelated sentences are joined together. We show that our approach is effective in a variety of other fields and topics, from the study of phase transitions and of frustration-induced disorder in statistical physics, to image classification.
Detecting and quantifying correlations in generic real-world data is a complex task since features live in high dimensional spaces and have interactions of arbitrary range and nature. For example, in an image, the meaning of a scene can emerge from the simultaneous presence of two objects that are physically far away. Moreover, semantic correlations are often associated with collective effects, such as the simultaneous presence of three or more specific objects. A prominent example where non-trivial correlations emerge is the feature space of deep neural networks. Those spaces typically lack a well-defined order, and neurons that are close in the architecture are not necessarily coding related information. However, the activations are known to be contained in a manifold of relatively low dimensionGoldt et al. [2020], and the existence of this manifold is believed to be an important ingredient in the learning process.
The main idea of this work is to use the intrinsic dimension (ID) of a dataset to detect and quantify correlations. If, for example, in a long text correlations between the first and the second part of the text vanish, the ID of the whole text should be approximately equal to the sum of the ID of the two parts. More precisely, in data points including features describing independent information, the ID should scale linearly with the number of features $N$, namely $ID=\gamma N$ for some positive constant $\gamma$. If one would be able to estimate the ID as a function of text length (or as a function of the patch dimension, in the case of images) one would be able to detect the range of the correlation. Moreover, the difference between the sum of the IDs of two representations and the ID of a dataset composed by the concatenation of these representations provides a quantitative measure of the strength of the correlationsBasile et al. [2025]: for perfectly correlated representations the ID of the concatenation will be equal to the ID of the single representations.
However, most of the available algorithms to perform ID estimation meet serious problems if the intrinsic dimension of the manifold is of the order of hundreds or moreCamastra and Staiano [2016]. In particular, some of them fail to provide an estimate of the ID which scales linearly with the number of features even when those are generated independently. Indeed, methods relying on the statistics of local neighborhoods systematically underestimate the ID because the number of samples required to sample a dataset grows exponentially in the IDEckmann and Ruelle [1992]. This problem is a manifestation of the so-called curse of dimensionalityBishop [2016] and affects to different extent all the ID estimators we are aware of, even if recently some progress has been madeErba et al. [2019].
@@
In order to estimate the ID in very large feature spaces, we here renounce to develop a general-purpose ID estimator, and we focus on features that are binary or can be approximately considered as binary. Concretely, we developed an ID estimator for binary data, called BID (from Binary Intrinsic Dimension) which, as we show below, can cope numerically with extremely large dimensionalities. Binary variables are the basis of all digital computation but are also of fundamental importance in both theoretical and applied science. The so-called spin systems, in which a spin is a binary degree of freedom, provide standard models to understand phases of matter, such as spin glassesBinder and Young [1986], and spin liquidsSavary and Balents [2016], and the corresponding phase transitions, like spontaneous orderingKardar [2007]. Binary variables are also used to model social dynamicsCastellano et al. [2009], biological neuronsHopfield [1982] and learning processesEngel and Broeck [2001], Barbier et al. [2025], Marino and Ricci-Tersenghi [2024]. In artificial neural networks, the activations are typically real numbers, but many efforts focus on finding more efficient representations: the so-called quantization methodsHubara et al. [2018], Guo [2018] aim to represent activations or weights using small integer numbers or binary variables, in order to drastically save memory, energy, and computing time, while preserving performance. Activations and weights can be transformed in binary variables by the sign functionHubara et al. [2016], Wang et al. [2023]. Binarization has been shown to preserve the scalar product between the activations and weightsAnderson and Berg [2018]. In Supp. Inf. we show that binarization also approximately preserves the local neighbourhoods computed using the full-precision activations inside a Large Language Model (LLM). This suggests that binarizing in high dimensions is an operation that retains relevant information about the dataGlielmo et al. [2022a].
In the following, we introduce an algorithm that can be used to estimate the binary intrinsic dimension (BID) of very large datasets of binary variables. We first show that it provides estimates that scale approximately linearly with system size when the system can be split into subsystems that are statistically independent, and reliably estimates the intrinsic dimension for bit streams of length up to $L\sim O(10^{6})$, even using only $O(1000)$ samples. We benchmark the algorithm on model systems in which the binary variables interact with an explicit energy function, which allows modulating their correlations. We show that the dependence of the BID on system size allows for characterizing the nature of correlations between the variables and spotting the transitions between ordered and disordered phases. We finally apply this approach to the study of the internal representations of images and text in large neural networks.
I Binomial estimator
@@
where $d_{0}$ and $d_{1}$ are variational parameters. In Supp. Inf. we show a Maximum Likelihood Estimation approach to infer a scale-dependent BID that supports our assumption of linear dependence (3).
@@
where $r^{*}$ is a meta-parameter that allows constraining the fit to small distances (see Supp. Inf.). As a representative measure of the BID of a system, we consider the limit for small distances of $d(r)$, which qualitatively corresponds to the local number of degrees of freedom. This definition is consistent with the idea that, if the data are defined by real variables, the ID should be equal to the dimension of the hyperplane that is tangent to the data manifold, which can be estimated considering only small distances. According to Eq. (3), we simply have $BID\equiv d_{0}$. The computational complexity of computing our observable $P_{emp}(r)$ scales linearly with the number of features and quadratically with the number of samples, since it involves computing all the Hamming distances between each pair of points.
Eq. (4) allows us to quantify rigorously how well our model approximates the observed data, $P_{emp}(r)$. In order to assign an uncertainty to the value obtained by the optimization, one can perform Bayesian inference with a Markov Chain Monte Carlo (MCMC) on the parameter space of model (2), given the observations. In Supp. Inf. we show that both the minimization of Eq. (4) and MCMC Bayesian inference with a uniform prior provide similar results. MCMC allows estimating the posterior distribution of two parameters of the model, and, as a consequence their mean and variances.
II Results
@@
We first show that our algorithm is able to compute intrinsic dimensions that scale linearly with size when the input data consists of statistically independent subsystems. We generated 2500 bit streams by concatenating $B$ blocks of $N=10^{4}$ bits, where each block is generated by independently harvesting spins from a Boltzmann distribution at $T=2$, with an energy function $E(\boldsymbol{\sigma})=\sum_{i=1}^{N=1}\sigma_{i}\sigma_{i+1}$ with periodic boundary conditions ($\sigma_{N+1}\equiv\sigma_{1}$). Table 1 shows that the BID per bit is approximately constant from $10^{4}$ bits to $10^{6}$ bits, consistent with the system being constructed by concatenating independent blocks. Remarkably, the approach allows estimating intrinsic dimensions of order $10^{5}$ using only $2500$ data points. This is empirical evidence that the estimator proposed in this work is not affected by the curse of dimensionality, and can be used in extremely high dimensional spaces. For a detailed analysis of the dependence on the number of samples see Sec. A.2 of Supp. Inf.
@@
Figure 1: The BID of spin systems as models of interacting bit arrays. Upper row: BID per spin as a function of the total number of spins $N$, for several temperatures $T$. Central row: Thermal dependence of the BID normalized per spin. $L$ stands for the lattice width (or height). Bottom row: Model validation at $L=100$ and different temperatures. (A,D,G): ferromagnetic Ising model on the square lattice, (B,E,H): the Potts model on the square lattice with $q=8$ states, (C,F,I): the antiferromagnetic Ising model on the triangular lattice. On vertical dashed lines the exact results in the thermodynamic limit for (D) the critical temperature, $T_{c}=2/\log(1+\sqrt{2})$ and (E) the transition temperature $T^{*}=1/\ln{(1+\sqrt{q})}\approx 0.745$.
@@
Next, we show that the BID is capable of capturing correlations between bits even when those span very large distances. We consider spins harvested from the Boltzmann distribution of the two-dimensional ferromagnetic Ising model on the square lattice (see Supp. Inf. for details). At high temperatures, spin-spin correlations are short-range, namely, they decay exponentially fast with the lattice distance. The characteristic scale of these correlations, the correlation length $\xi(T)$, is much smaller than the system size $L=\sqrt{N}$, implying that the BID must be asymptotically proportional to the number of bits. This is observed in the curve corresponding to $T=4.0$ in Fig. 1A. The value of the BID per bit is close to 1, reflecting the presence of small correlations. The same behavior is observed at $T=3$, with a relative BID that is smaller, consistently with the fact that correlations become more important. At low temperatures, the system is globally ordered (namely the majority of bits are either 00 or $1$) but the fluctuations are also short-range correlatedBinney et al. [1992]. Thus the BID is again proportional to the number of bits, but its value is much lower, reflecting the presence of a dominant strong correlation between the variables "forcing" most of the bits to align. Near the critical temperature ($T=2.3$) the relative BID does not reach a plateau, indicating that the correlation length is comparable with the size of the system. This is a well-known phenomenon observed in spin systems at criticalityBinney et al. [1992], which is captured by the BID through the sub-linear scaling with the number of bits. Fig. 1D shows the BID as a function of temperature for the 2D ferromagnetic Ising model. Away from criticality, the BID per spin tends smoothly to zero or one as the temperature goes to zero or infinity, respectively. The critical temperature is easily recognized in the graph since, consistently with ref. Mendes-Santos et al. [2021], we observe a local minimum of the BID at that temperature.
@@
The second system we studied is a $q$-state Potts model, with $q=8$ (see Supp. Inf. for details). In order to calculate its BID we construct a binary system by writing the state (or color) index $q$ in a binary representation. The correlation length in this model has a maximum at the transition temperature but remains finite even in the thermodynamic limitBuddenoir and Wallon [1993]. This behavior is correctly captured by the BID scaling being asymptotically linear for every temperature (Fig. 1B), even close to the transition temperature $T^{*}$. The BID shows an abrupt change around $T^{*}$, consistent with the underlying first-order thermal phase transition.
@@
Finally, we considered the Ising model on the triangular lattice with antiferromagnetic first-neighbor interactions. Due to geometrical frustration, this model does not have a thermal phase transition but presents a smooth crossover towards a ground state that has an entropy proportional to the number of spins in the systemWannier [1950] and moreover presents a power-law decay of correlationsBlöte and Nightingale [1993]. Fig. 1C captures the long-range correlations of the ground-state through the sublinear scaling of the BID per bit, only present at small temperatures. Decreasing the temperature, the BID per spin tends smoothly to a non-zero value for all system sizes studied (Fig. 1F). The bottom panels in Fig. 1 illustrate the viability of the model defined in eq. Eqs 2 and 3 to describe the observed histograms of the distances between bit streams in all the conditions considered in the tests: the black lines, corresponding to the model of Eq. 2, fit pretty accurately the observations in all the conditions, even if at different temperatures the observations are concentrated at very different distances.
@@
Figure 2: The BID of image-crop representations in Resnet18. Upper row of images: example of crops for a sample of ImageNet. On top of each image, we report the size of the crop. All images are resized to $224\times 224$ resolution after cropping, thus the rightmost image is the full image. Note moreover that the crop is performed only along the spatial dimensions, keeping the three channels of the images. Lower row of images: a set of images constructed from random patches of size $28\times 28$ from ImageNet samples. A, B, C, D panels: BID of image representations as a function of the number of pixels in the cropped image, $N_{crop}$, where we normalized the latter by dividing by the total number of pixels in the full image, $N_{tot}=224^{2}$. For each panel, $l$ is the layer index and $L_{R}$ is the total number of layers. For further details see Supp. Inf.
These benchmark examples show that BID is able to characterize the collective behavior of a large number of bits. Analogously, in computer vision, the collective behavior of a large number of pixels gives rise to patterns, which encode the meaning of an image. We computed the BID of internal representations of a convolutional image classifier, constructing binary data representations simply using the sign of the real-valued activations (see Supp. Inf. for details). The BID behaves as a proxy of semantic content, as we show through the following size-scaling experiment. The first row of Fig. 2 shows an example of different crops for one element of ImageNet, where each crop is resized to the fixed full-image resolution. Fig. 2 shows how the BID of the representations inside Resnet18 depends on the initial crop size, for different layers. For small crop sizes the BID increases monotonically, since on average the semantic content related to the classification task is being added. For large enough crop sizes, the BID saturates, indicating that all the relevant content of the image was already present in previous sizes. Later layers saturate earlier since they carry information that is only relevant for the classification task and on average adding more context does not make significant changes. Indeed, if we consider a sufficiently small crop of a generic image, it is not possible to assign a label to it, since it does not carry enough information to make a discrimination. Instead, semantic information emerges for big enough crop sizes but saturates when the crop is so large that the extra pixels do not bring new information. In Fig. 2 we also show the same analysis performed on images constructed by an ensemble of random patches of the same data set. In this case, the semantic content is lost, even if the local structure in the images is preserved. Consequently, the BID does not plateau, since large images correspond to a collection of independent crops, each bringing the addition of independent information.
@@
Figure 3: The BID of text representations in large language models. Upper boxes: an example of a Wikitext sentence truncated at different token lengths, $T$. Central panel: BID per bit calculated using all token representations in the sequence as a function of the number of tokens, for Pythia410m (see Supp. Inf. for details). $l=0$ stands for the embedding layer, $l=24$ corresponds to the last transformer layer before the head. "Concat" stands for phrases constructed concatenating 150 tokens from two randomly selected data samples. The dashed line corresponds to the semantic boundary between the two phrases.
@@
As a last example, we show that the BID can be used to detect correlations in the feature maps learned by large transformer-based architectures trained on a natural language processing (NLP) task. The ID has been used to analyze the properties of deep autoregressive transformers for protein sequence analysis and image analysisValeriani et al. [2023] and more recently for language Cheng et al. [2023, 2025]. However, to make the ID estimation possible in Ref. Valeriani et al. [2023] the representations were averaged in the sequence dimension, mixing information associated with the different tokens. In Refs. Cheng et al. [2023, 2025], the analyses are performed only on the representation of the last token. We here use our approach to estimate the binary intrinsic dimension of the representation of whole sentences, probing in this manner the collective properties of all the neurons of a layer. We consider paragraphs from Wikitext consisting of more than $300$ tokens, and we study how the BID scales with text length, binarizing activations through the application of the sign function. The initial representation (layer 0, Fig. 3, left panel) has short-range correlations, signaled by the BID per bit slowly drifting up as a function of the number of tokens (Fig. 3). Token embeddings are correlated because they must be compliant with the grammar structure of natural language, in which words are often used in groups. For long sentences, the BID scales approximately linearly with the sentence length: grammar-related correlations become less relevant in parts of a sentence separated by hundreds of words. Instead, the BID per bit of the last-layer representation learned by the model displays a clear power-law decay as a function of the number of tokens, similar to the one observed in scale-free critical spin systems (see Fig. 1A). Since the semantic content in text is a collective property of the whole sentence, the network to capture this property develops a long-range correlated representation: qualitatively the process of understanding the meaning of a long sentence can be described as the process of finding meaningful correlations in the sequence of tokens. To validate that we are indeed capturing the semantic correlations in text through the BID scaling we repeat the experiment by concatenating 150 tokens from two randomly chosen sentences from the same corpus. We observe that the power law decrease stops precisely at the boundary between phrases, and it is followed by an approximate plateau, consistent with the two halves of the paragraph containing independent information. In the last part of the plot, the BID per bit starts decreasing again, as a result of the semantic correlations of the second half of the sentence. In Supp. Inf. we show that these results hold also for changing both corpus and network.
@@
III Comparison with other methods
Figure 4: Comparison between BID and other real-space estimators. Left panel: thermal dependence of the intrinsic dimension for a small $30\times 30$ ferromagnetic Ising model. Right panel: scaling of the intrinsic dimension normalized per bit at temperature $T=1.9$. Number of samples: 5000.

### v3

@@ Abstract
Introduction
Correlations among the features defining a data point force a data set to live in a manifold $\mathcal{M}$ whose geometry encodes the correlation structure of data. The most fundamental property of $\mathcal{M}$ is its intrinsic dimensionBennett [1969] (ID) which is qualitatively defined by the number of independent variables needed to parameterize it without significant information loss. If one was able to estimate the intrinsic dimension in a generic big data scenario, one would be able to quantify the strength of correlations in real-world datasets, like a corpus of sentences or a large set of measurements from a camera. However, real-world data are represented in feature spaces that are outrageously large, and most of the available methods significantly struggle if the intrinsic dimension of the manifold is of the order of hundreds or more Camastra and Staiano [2016]. Detecting and quantifying semantic correlations in generic real world data is a complex task since features live in high dimensional spaces and have interactions of arbitrary range and nature. For example, in an image, the meaning of a scene can emerge from the simultaneous presence of two objects that are physically far away. Moreover, semantic correlations are often associated with collective effects, such as the simultaneous presence of three or more specific objects. A prominent example where non-trivial correlations emerge is the feature space of deep neural networks. Those spaces typically lack a well-defined order, and neurons that are close in the architecture are not necessarily coding related information. However, the activations are known to be contained in a manifold of relatively low dimensionGoldt et al. [2020], and the existence of this manifold is believed to be an important ingredient in the learning process. The main idea of this work is to use the intrinsic dimension (ID) of a dataset to detect and quantify correlations. If, for example, in a long text correlation between the first and the second part of the text vanish, the ID of the whole text should be approximately equal to the sum of the ID of the two parts. More precisely, in data points including features describing independent information, the ID should scale linearly with the number of features N, namely $ID=\gamma N$ for some positive constant $\gamma$. If one were able to estimate the ID as a function of text length (or as a function of the patch dimension, in the case of images) one would be able to detect the range of the correlation. Moreover, the difference between the sum of the IDs of two representations and the ID of a dataset composed by the concatenation of these representations provides a quantitative measure of the strength of the correlationsBasile et al. [2025]: for perfectly correlated representations the ID of the concatenation will be equal to the ID of the single representations. However, most of the available algorithms to perform ID estimation meet serious problems if the intrinsic dimension of the manifold is of the order of hundreds or moreCamastra and Staiano [2016]. In particular, some of them fail to provide an estimate of the ID which scales linearly with the number of features even when those are generated independently. Indeed, methods relying on the statistics of local neighborhoods systematically underestimate the ID because the number of samples required to sample a dataset grows exponentially in the IDEckmann and Ruelle [1992]. This problem is a manifestation of the so-called curse of dimensionalityBishop [2016] and affects to different extent all the ID estimators we are aware of, even if recently some progress has been madeErba et al. [2019].
@@ Introduction
Binary variables are the basis of all digital computation but are also of fundamental importance in both theoretical and applied science. The so-called spin systems, in which a spin is a binary degree of freedom, provide standard models to understand phases of matter, such as spin glassesBinder and Young [1986], and spin liquidsSavary and Balents [2016], and the corresponding phase transitions, like spontaneous orderingKardar [2007]. Binary variables are also used to model social dynamicsCastellano et al. [2009], biological neuronsHopfield [1982] and learning processesEngel and Broeck [2001], Barbier et al. [2025], Marino and Ricci-Tersenghi [2024]. In artificial neural networks, the activations are typically real numbers, but many efforts focus on finding more efficient representations: the so-called quantization methodsHubara et al. [2018], Guo [2018] aim to represent activations or weights using small integer numbers or binary variables, in order to drastically save memory, energy, and computing time, while preserving performance. Activations and weights can be transformed in binary variables by the sign functionHubara et al. [2016], Wang et al. [2023], which has been shown to preserve the scalar product between themAnderson and Berg [2018]. Here, we introduce an intrinsic dimension estimator that generalizes the concept of the intrinsic dimension to the case of binary variables. We present an algorithm that can be used to estimate the binary intrinsic dimension (BID) of very large datasets of binary variables. We first show that it provides estimates that scale approximately linearly with system size when the system can be split into subsystems that are statistically independent. Concretely, we test the algorithm up to one million spins, using only 2500 samples. Then we benchmark the algorithm on model systems in which the binary variables interact with an explicit energy function, which allows modulating their correlations. We show that the dependence of the BID on system size allows for characterizing the nature of correlations between the variables and spotting phase transitions. In Supplementary Note 4 We show that binarization also approximately preserves the local neighbourhoods computed using the full-precision activations inside a Large Language Model (LLM). This suggests that binarizing in high dimensions is an operation that retains relevant information about the dataGlielmo et al. [2022a]. We finally apply this approach to the study of the internal representations of images and text in large neural networks. This result paves the way to a plethora of applications, allowing to address questions which have thus far been, to the best of our knowledge, intractable. Our approach is effective in a variety of fields and topics, from the study of phase transitions and of frustration-induced disorder in statistical physics, to image classification.
Results
Theoretical model
@@ Results
where $d_{0}$ and $d_{1}$ are variational parameters. In Methods, section ‘Maximum Likelihood Estimation of a scale-dependent BID’ we show a Maximum Likelihood Estimation approach to infer a scale-dependent BID that supports our assumption of linear dependence (3).
@@
where $r^{*}$ is a meta-parameter that allows constraining the fit to small distances (see Supplementary Note 1). As a representative measure of the BID of a system, we consider the limit for small distances of $d(r)$, which qualitatively corresponds to the local number of degrees of freedom. This definition is consistent with the idea that, if the data are defined by real variables, the ID should be equal to the dimension of the hyperplane that is tangent to the data manifold, which can be estimated considering only small distances. According to Eq. (3), we simply have $BID\equiv d_{0}$. The computational complexity of computing our observable $P_{emp}(r)$ scales linearly with the number of features and quadratically with the number of samples, since it involves computing all the Hamming distances between each pair of points.
Eq. (4) allows us to quantify rigorously how well our model approximates the observed data, $P_{emp}(r)$. In order to assign an uncertainty to the value obtained by the optimization, one can perform Bayesian inference with a Markov Chain Monte Carlo (MCMC) on the parameter space of model (2), given the observations. In Methods, section ‘Bayesian inference of the BID by Markov Chain Monte Carlo’ we show that both the minimization of Eq. (4) and MCMC Bayesian inference with a uniform prior provide similar results. MCMC allows estimating the posterior distribution of two parameters of the model, and, as a consequence their mean and variances.
Numerical experiments
@@
We first show that our algorithm is able to compute intrinsic dimensions that scale linearly with size when the input data consists of statistically independent subsystems. We generated 2500 bit streams by concatenating $B$ blocks of $N=10^{4}$ bits, where each block is generated by independently harvesting spins from a Boltzmann distribution at $T=2$, with an energy function $E(\boldsymbol{\sigma})=\sum_{i=1}^{N=1}\sigma_{i}\sigma_{i+1}$ with periodic boundary conditions ($\sigma_{N+1}\equiv\sigma_{1}$). Table 1 shows that the BID per bit is approximately constant from $10^{4}$ bits to $10^{6}$ bits, consistent with the system being constructed by concatenating independent blocks. Remarkably, the approach allows estimating intrinsic dimensions of order $10^{5}$ using only $2500$ data points. This is empirical evidence that the estimator proposed in this work is not affected by the curse of dimensionality, and can be used in extremely high dimensional spaces. For a detailed analysis of the dependence on the number of samples see Supplementary Note 2.
@@
Figure 1: The Binary Intrinsic Dimension (BID) of spin systems as models of interacting bit arrays. Upper row: BID per spin as a function of the total number of spins, $N$, for several temperatures $T$. Central row: Thermal dependence of the BID normalized per spin. $L$ stands for the lattice width (or height). Bottom row: Model validation at $L=100$ and different temperatures. In colored crosses, the empirical probability of Hamming distances $P_{emp}(r)$. In solid black line, our model fit, Eq. (2). (a,d,g): ferromagnetic Ising model on the square lattice, (b,e,h): the Potts model on the square lattice with $q=8$ states, (c,f,i): the antiferromagnetic Ising model on the triangular lattice. With vertical dashed lines we show the exact results in the thermodynamic limit for the critical temperature, $T_{c}=2/\log(1+\sqrt{2})$ in panel d) and the transition temperature $T^{*}=1/\ln{(1+\sqrt{q})}\approx 0.745$ in panel e).
@@
Next, we show that the BID is capable of capturing correlations between bits even when those span very large distances. We consider spins harvested from the Boltzmann distribution of the two-dimensional ferromagnetic Ising model on the square lattice (see Methods, section ‘Spin systems’ for details). At high temperatures, spin-spin correlations are short-range, namely, they decay exponentially fast with the lattice distance. The characteristic scale of these correlations, the correlation length $\xi(T)$, is much smaller than the system size $L=\sqrt{N}$, implying that the BID must be asymptotically proportional to the number of bits. This is observed in the curve corresponding to $T=4.0$ in Fig. 1a). The value of the BID per bit is close to 1, reflecting the presence of small correlations. The same behavior is observed at $T=3$, with a relative BID that is smaller, consistently with the fact that correlations become more important. At low temperatures, the system is globally ordered (namely the majority of bits are either 00 or $1$) but the fluctuations are also short-range correlatedBinney et al. [1992]. Thus the BID is again proportional to the number of bits, but its value is much lower, reflecting the presence of a dominant strong correlation between the variables "forcing" most of the bits to align. Near the critical temperature ($T=2.3$) the relative BID does not reach a plateau, indicating that the correlation length is comparable with the size of the system. This is a well-known phenomenon observed in spin systems at criticalityBinney et al. [1992], which is captured by the BID through the sub-linear scaling with the number of bits. Fig. 1d) shows the BID as a function of temperature for the 2D ferromagnetic Ising model. Away from criticality, the BID per spin tends smoothly to zero or one as the temperature goes to zero or infinity, respectively. The critical temperature is easily recognized in the graph since, consistently with ref. Mendes-Santos et al. [2021], we observe a local minimum of the BID at that temperature.
@@
The second system we studied is a $q$-state Potts model, with $q=8$ (see Methods, section ‘Spin systems’ for details). In order to calculate its BID we construct a binary system by writing the state (or color) index $q$ in a binary representation. The correlation length in this model has a maximum at the transition temperature but remains finite even in the thermodynamic limitBuddenoir and Wallon [1993]. This behavior is correctly captured by the BID scaling being asymptotically linear for every temperature (Fig. 1b), even close to the transition temperature $T^{*}$. The BID shows an abrupt change around $T^{*}$, consistent with the underlying first-order thermal phase transition.
@@
Finally, we considered the Ising model on the triangular lattice with antiferromagnetic first-neighbor interactions. Due to geometrical frustration, this model does not have a thermal phase transition but presents a smooth crossover towards a ground state that has an entropy proportional to the number of spins in the systemWannier [1950] and moreover presents a power-law decay of correlationsBlöte and Nightingale [1993]. Fig. 1c) captures the long-range correlations of the ground-state through the sublinear scaling of the BID per bit, only present at small temperatures. Decreasing the temperature, the BID per spin tends smoothly to a non-zero value for all system sizes studied (Fig. 1f).
The bottom panels in Fig. 1 illustrate the viability of the model defined in eq. Eqs 2 and 3 to describe the observed histograms of the distances between bit streams in all the conditions considered in the tests: the black lines, corresponding to the model of Eq. 2, fit pretty accurately the observations in all the conditions, even if at different temperatures the observations are concentrated at very different distances.
@@
Figure 2: The Binary Intrinsic Dimension (BID) of image-crop representations in Resnet18. a) Example of crops for a sample of ImageNet. On top of each image, we report the size of the crop. All images are resized to $224\times 224$ resolution after cropping, thus the rightmost image is the full image. Note moreover that the crop is performed only along the spatial dimensions, keeping the three channels of the images. b): a set of images constructed from random patches of size $28\times 28$ from ImageNet samples. c), d), e), f) panels: BID of image representations as a function of the number of pixels in the cropped image, $N_{crop}$, where we normalized the latter by dividing by the total number of pixels in the full image, $N_{tot}=224^{2}$. For each panel, $l$ is the layer index and $L_{R}$ is the total number of layers. For further details see Methods, section ‘The BID of image representations’.
@@
Figure 3: The Binary Intrinsic Dimension (BID) of text representations in large language models. Upper boxes: an example of a Wikitext sentence truncated at different token lengths, $T$. Central panel: BID per bit calculated using all token representations in the sequence as a function of the number of tokens, for Pythia410m (see Methods, section ‘The BID of text representations’ for details). $l=0$ stands for the embedding layer, $l=24$ corresponds to the last transformer layer before the head. "Concat" stands for phrases constructed concatenating 150 tokens from two randomly selected data samples. The dashed line corresponds to the semantic boundary between the two phrases.
These benchmark examples show that BID is able to characterize the collective behavior of a large number of bits. Analogously, in computer vision, the collective behavior of a large number of pixels gives rise to patterns, which encode the meaning of an image. We computed the BID of internal representations of a convolutional image classifier, constructing binary data representations simply using the sign of the real-valued activations (see Methods, section ‘The BID of image representations’ for details). The BID behaves as a proxy of semantic content, as we show through the following size-scaling experiment. Fig. 2a) shows an example of different crops for one element of ImageNet, where each crop is resized to the fixed full-image resolution. Fig. 2c)-f) shows how the BID of the representations inside Resnet18 depends on the initial crop size, for different layers. For small crop sizes the BID increases monotonically, since on average the semantic content related to the classification task is being added. For large enough crop sizes, the BID saturates, indicating that all the relevant content of the image was already present in previous sizes. Later layers saturate earlier since they carry information that is only relevant for the classification task and on average adding more context does not make significant changes. Indeed, if we consider a sufficiently small crop of a generic image, it is not possible to assign a label to it, since it does not carry enough information to make a discrimination. Instead, semantic information emerges for big enough crop sizes but saturates when the crop is so large that the extra pixels do not bring new information. In Fig. 2c-f) we also show the same analysis performed on images constructed by an ensemble of random patches of the same data set. In this case, the semantic content is lost, even if the local structure in the images is preserved. Consequently, the BID does not plateau, since large images correspond to a collection of independent crops, each bringing the addition of independent information.
@@
As a last example, we show that the BID can be used to detect correlations in the feature maps learned by large transformer-based architectures trained on a natural language processing (NLP) task. The ID has been used to analyze the properties of deep autoregressive transformers for protein sequence analysis and image analysisValeriani et al. [2023] and more recently for language Cheng et al. [2023, 2025]. However, to make the ID estimation possible in Ref. Valeriani et al. [2023] the representations were averaged in the sequence dimension, mixing information associated with the different tokens. In Refs. Cheng et al. [2023, 2025], the analyses are performed only on the representation of the last token. We here use our approach to estimate the binary intrinsic dimension of the representation of whole sentences, probing in this manner the collective properties of all the neurons of a layer. We consider paragraphs from Wikitext consisting of more than $300$ tokens, and we study how the BID scales with text length, binarizing activations through the application of the sign function. The initial representation (layer 0, Fig. 3, left panel) has short-range correlations, signaled by the BID per bit slowly drifting up as a function of the number of tokens (Fig. 3). Token embeddings are correlated because they must be compliant with the grammar structure of natural language, in which words are often used in groups. For long sentences, the BID scales approximately linearly with the sentence length: grammar-related correlations become less relevant in parts of a sentence separated by hundreds of words. Instead, the BID per bit of the last-layer representation learned by the model displays a clear power-law decay as a function of the number of tokens, similar to the one observed in scale-free critical spin systems (see Fig. 1a). Since the semantic content in text is a collective property of the whole sentence, the network to capture this property develops a long-range correlated representation: qualitatively the process of understanding the meaning of a long sentence can be described as the process of finding meaningful correlations in the sequence of tokens. To validate that we are indeed capturing the semantic correlations in text through the BID scaling we repeat the experiment by concatenating 150 tokens from two randomly chosen sentences from the same corpus. We observe that the power law decrease stops precisely at the boundary between phrases, and it is followed by an approximate plateau, consistent with the two halves of the paragraph containing independent information. In the last part of the plot, the BID per bit starts decreasing again, as a result of the semantic correlations of the second half of the sentence. In Methods, section ‘The BID of text representations’ we show that these results hold also for changing both corpus and network.
@@
Comparison with other methods
Figure 4: Comparison between Binary Intrinsic Dimension (BID) and other real-space estimators. Panel a): Thermal dependence of the intrinsic dimension for a small $30\times 30$ ferromagnetic Ising model. Panel b) scaling of the intrinsic dimension normalized per bit at temperature $T=1.9$. Number of samples: 5000. GRIDE stands for the Generalized Ratios Intrinsic Dimension EstimatorDenti et al. [2022]. FCI stands for the Full Correlation Integral estimatorErba et al. [2019]. CD stands for Correlation IntegralGrassberger and Procaccia [1983]. MLE stands for the Maximum Likelihood Estimator of Ref. Levina and Bickel [2004]. I3D stands for the Intrinsic Dimension Estimator for Discrete Datasets of Ref. Macocco et al. [2023].